# Sample-Specific Noise Injection For Diffusion-Based Adversarial Purification

### v1

@@ 6 Limitation
Maximum level $t^{\rm S}$ for EPS. We use EPS to replace the single reference noise level with an integrated approach. Still, the maximum level $t^{\rm S}$ defining the upper bound of the expectation range is shared across samples, which may not be the optimal choice though. We leave the exploration of more sample-sensitive maximum levels to future studies.
The design of reweighting functions. The proposed reweighting functions (i.e., the linear and non-linear reweighting functions) may not be the optimal ones for SSNI. However, designing an effective reweighting function is an open question, and we leave it as future work.
Extra computational cost. The integration of an extra reweighting process will inevitably bring some extra cost. Luckily, we find that this reweighting process is lightweight, making SSNI computationally feasible compared to existing DBP methods (see Section 5.6).
@@ Acknowledgements
JCZ is supported by the Melbourne Research Scholarship and would like to thank Shuhai Zhang and Huanran Chen for productive discussions. ZSY is supported by the Australian Research Council (ARC) with grant number DE240101089. FL is supported by the ARC with grant number DE240101089, LP240100101, DP230101540 and the NSF&CSIRO Responsible AI program with grant number 2303037. This research was supported by The University of Melbourne’s Research Computing Services and the Petascale Campus Initiative.
@@ Appendix A Detailed Related Work
Adversarial attack. Adversarial examples (AEs) have emerged as a critical security concern in the development of AI systems in recent years (Szegedy et al., 2014; Goodfellow et al., 2015). These examples are typically generated by introducing imperceptible perturbations to clean inputs, which can cause a classifier to make incorrect predictions with high confidence. The methods used to generate such examples are known as adversarial attacks. One of the earliest attack methods, the fast gradient sign method (FGSM), which perturbs clean data in the direction of the gradient of the loss function (Goodfellow et al., 2015). Building on this idea, Madry et al. (2018) propose the projected gradient descent (PGD) attack, which applies iterative gradient-based updates with random initialization. AutoAttack (AA) (Croce & Hein, 2020) integrate multiple attack strategies into a single ensemble, making it a widely adopted benchmark for evaluating adversarial robustness. To address defenses that apply randomized input transformations, Athalye et al. (2018b) introduce the expectation over transformation (EOT) framework for computing more accurate gradients. Additionally, Athalye et al. (2018a) propose the backward pass differentiable approximation (BPDA), which approximates gradients using identity mappings to circumvent gradient obfuscation defenses. According to Lee & Kim (2023), the combination of PGD and EOT, i.e., PGD+EOT, is currently considered the most effective attack strategy against DBP methods. More recently, adversarial attacks that specifically designed for DBP methods are proposed. For example, Xue et al. (2023) propose diffusion-based projected gradient descent (Diff-PGD), which leverages an off-the-shelf diffusion model to guide perturbation optimization, enabling the generation of more stealthy AEs. Chen et al. (2024b) propose DiffAttack, which crafts perturbations in the latent space of diffusion models, rather than directly in pixel space, enabling the generation of human-insensitive yet semantically meaningful AEs through content-preserving structures.
@@
Adversarial defense. To counter the threat posed by adversarial attacks, numerous defense strategies have been developed, including adversarial detection (AD), adversarial training (AT), and adversarial purification (AP). (1) AD: AD is the most lightweight approach to defending against adversarial attacks is to detect and remove AEs from the input data. Earlier studies typically trained detectors tailored to specific classifiers or attack types, often overlooking the underlying data distribution, which limits their generalization to unseen attacks (Ma et al., 2018; Lee et al., 2018; Raghuram et al., 2021; Pang et al., 2022). Recently, statistical adversarial data detection has attracted growing attention for its ability to address this limitation. For instance, Gao et al. (2021) show that the maximum mean discrepancy (MMD) is sensitive to adversarial perturbations, and use distributional discrepancies between AEs and CEs to effectively identify and filter out AEs, even under previously unseen attacks. Building on this insight, Zhang et al. (2023) introduce a novel statistic called the expected perturbation score (EPS), which quantifies the average score of a sample after applying multiple perturbations. They then propose an EPS-based variant of MMD to capture the distributional differences between clean and adversarial examples more effectively. (2) AT: Vanilla AT (Madry et al., 2018) directly generates adversarial examples during training, encouraging the model to learn their underlying distribution. Beyond vanilla AT, various extensions have been proposed to improve its effectiveness. For instance, Zhang et al. (2019) introduce a surrogate loss optimized based on theoretical robustness bounds. Similarly, Wang et al. (2020) examine the role of misclassified examples in shaping model robustness and enhance adversarial risk via regularization techniques. From a reweighting perspective, Zhang et al. (2021) propose geometry-aware instance-reweighted AT (GAIRAT), which adjusts sample weights based on their distance to the decision boundary. Building on this, Wang et al. (2021) utilize probabilistic margins to reweight AEs in a continuous and path-independent manner. More recently, Zhang et al. (2024) suggest pixel-wise reweighting of AEs to explicitly direct attention toward critical image regions. (3) AP: AP typically employs generative models to transform AEs back into their clean counterparts before classification (Liao et al., 2018; Samangouei et al., 2018; Song et al., 2018; Naseer et al., 2020; Zhang et al., 2025). Within this context, diffusion-based purification (DBP) methods have emerged as a promising framework, exploiting the inherent denoising nature of diffusion models to filter out adversarial noise (Nie et al., 2022; Wang et al., 2022; Xiao et al., 2023; Lee & Kim, 2023). DBP works with pre-trained diffusion model by first corrupting an adversarial input through the forward process and then iteratively reconstructing it via the reverse process. This projects the input back onto the clean data manifold while stripping away adversarial artifacts. Wang et al. (2022) introduce input guidance during the reverse diffusion process to ensure the purified outputs stay close to the inputs. Lee & Kim (2023) propose a fine-tuned gradual noise scheduling for multi-step purifications. Bai et al. (2024) improve the reverse diffusion process by incorporating contrastive objectives.
@@
Let ${\mathbf{y}}\in\Omega_{\rm SH}({\mathbf{x}})$. Then $\exists t^{*}\in[0,T_{\rm max}]$ such that ${\mathbf{y}}=R({\mathbf{x}}_{t^{*}})$. Define $f(z,t^{*})=t^{*}$ for all $z$ and $t^{*}$, then $t({\mathbf{x}})=f(||s_{\theta}({\mathbf{x}},t^{*})||,t^{*})=t^{*}$. Therefore, ${\mathbf{y}}=R({\mathbf{x}}_{t^{*}})=R({\mathbf{x}}_{t({\mathbf{x}})})\in\Phi_{\rm SI}({\mathbf{x}})$. This completes the proof of $\Omega_{\rm SH}({\mathbf{x}})\subseteq\Omega_{\rm SI}({\mathbf{x}})$. ∎
@@ C.1 DiffPure
Existing DBP methods generally follow the algorithm of DiffPure (Nie et al., 2022). DiffPure conducts evaluation on AutoAttack (Croce & Hein, 2020) and BPDA+EOT adaptive attack (Athalye et al., 2018a) to measure model robustness. DiffPure chooses optimal $t^{*}=100$ and $t^{*}=75$ on CIFAR-10 against threat models $\ell_{\infty}(\epsilon=8/255)$ and $\ell_{2}(\epsilon=0.5)$, respectively. It also tests on high-resolution dataset like ImageNet-1K with $t^{*}=150$ against threat models $\ell_{\infty}(\epsilon=4/255)$. Calculating exact full gradients of the defense system of DiffPure is impossible since one attack iteration requires 100 function calls (with $t^{*}=100$ and a step size of 1). DiffPure originally uses numerical SDE solver for calculating gradients. However, the adjoint method is insufficient to measure the model robustness since it relies on the performance of the underlying SDE solver (Zhuang et al., 2020; Lee & Kim, 2023; Chen et al., 2024a). Therefore, we apply surrogate process to efficiently compute gradients of direct back-propagation in our evaluation. To overcome memory constraint issue, we align the step size settings of denoising process in adversarial attack to 5 with the evaluation settings in (Lee & Kim, 2023) and keep the timestep $t^{*}$ consistent with DiffPure. For ImageNet-1K evaluation, we can only afford a maximum of 10 function calls for one attack iteration.
@@ C.2 GDMP
Lee & Kim (2023) evaluated GDMP with three types of guidance and concluded that No-Guidance provides the best robustness performance when using the surrogate process to compute the full gradient through direct backpropagation. In our evaluation, we incorporate the surrogate process with No-Guidance to evaluate GDMP. Since it is impossible to calculate the gradients of the full defense system, we use a surrogate process consisting of same number of purification steps but with larger step size in the attack (with 6 denoising steps and 10 denoising steps for PGD+EOT and BPDA+EOT attack, respectively). Notably, GDMP only uses one purification run with 45 forward steps to evaluate on ImageNet-1K, which we keep consistent with this setting.
@@ C.3 GNS
Lee & Kim (2023) emphasizes the importance of selecting optimal hyperparameters in DBP methods for achieving better robust performance. Hence, Lee & Kim (2023) proposed Gradual Noise-Scheduling (GNS) for multi-step purification, which is based on the idea of choosing the best hyperparameters for multiple purification steps. It is basically the same architecture as GDMP (no guidance), but with different purification steps, forward steps and denoising steps. Specifically, GNS sets different forward and reverse diffusion steps and gradually increases the noise level at each subsequent purification step. We just keep the same hyperparameter settings and also use an ensemble of 10 runs to evaluate the method.
@@ Appendix D Surrogate Process of Gradient Computation
The surrogate process is an efficient approach for computing approximate gradients through backpropagation, as proposed by (Lee & Kim, 2023). White-box adaptive attacks, such as PGD+EOT, involve an iterative optimization process that requires computing the exact full gradients of the entire system, result in high memory usage and increased computational time. DBP methods often include a diffusion model as an add-on purifier, which the model requires extensive function calls during reverse generative process. Hence, it is hard to compute the exact full gradient of DBP systems efficiently. The surrogate process takes advantage of the fact that, given a fixed total amount of noise, we can denoise it using fewer denoising steps (Song et al., 2021a), but the gradients obtained from the surrogate process slightly differ from the exact gradients. Instead of using the full denoising steps, we approximate the original denoising process with fewer function calls, which allows us to compute gradients by directly back-propagating through the forward and reverse processes.
There are other gradient computation methods such as adjoint method in DiffPure (Li et al., 2020; Nie et al., 2022). It leverages an underlying numerical SDE solver to solve the reverse-time SDE. The adjoint method can theoretically compute exact full gradient, but in practice, it relies on the performance of the numerical solver, which is insufficient to measure the model robustness (Zhuang et al., 2020; Lee & Kim, 2023; Chen et al., 2024a). Lee & Kim (2023) conducted a comprehensive evaluation of both gradient computation methods and concluded that utilizing the surrogate process for gradient computation poses a greater threat to model robustness. Hence, we use gradients obtained from a surrogate process in all our experiments.
@@ Appendix E Adaptive White-box PGD+EOT Attack for SSNI
0:  clean data-label pairs (${\mathbf{x}},y)$; purifier $f_{p}$; classifier $f_{c}$; a noise level $T$; a score network $s_{\theta}({\mathbf{x}},T)$; objective function ${\mathcal{L}}$; perturbation budget $\epsilon$; step size $\alpha$; PGD iterations $K$; EOT iterations $N$.
@@ Appendix F Adaptive White-box BPDA+EOT Attack for SSNI
0:  clean data-label pairs (${\mathbf{x}},y)$; purifier $f_{p}$; classifier $f_{c}$; approximation function $f_{\text{apx}}$; a noise level $T$; a score network $s_{\theta}({\mathbf{x}},T)$; objective function ${\mathcal{L}}$; perturbation budget $\epsilon$; step size $\alpha$; PGD iterations $K$; EOT iterations $N$.
@@ Appendix G Performance Evaluation of SSNI-L
WRN-28-10 Nie et al. (2022) 89.71$\pm$0.72 47.98$\pm$0.64
+ SSNI-L 92.97$\pm$0.42 46.35$\pm$0.72
Wang et al. (2022) 92.45$\pm$0.64 36.72$\pm$1.05
+ SSNI-L 93.62$\pm$0.49 36.59$\pm$1.29
Lee & Kim (2023) 90.1$\pm$0.18 56.05$\pm$1.11
+ SSNI-L 93.49$\pm$0.33 53.71$\pm$0.48
WRN-70-16 Nie et al. (2022) 90.89$\pm$1.13 52.15$\pm$2.30
+ SSNI-L 93.82$\pm$0.49 49.94$\pm$0.33
Wang et al. (2022) 93.10$\pm$0.51 43.55$\pm$0.58
+ SSNI-L 93.88$\pm$0.49 43.03$\pm$0.60
Lee & Kim (2023) 89.39$\pm$1.12 56.97$\pm$0.33
+ SSNI-L 92.64$\pm$0.40 52.86$\pm$0.46
@@
WRN-28-10 Nie et al. (2022) 91.80$\pm$0.84 82.81$\pm$0.97
+ SSNI-L 93.82$\pm$0.37 81.12$\pm$0.80
Wang et al. (2022) 92.45$\pm$0.64 82.29$\pm$0.82
+ SSNI-L 93.62$\pm$0.49 80.66$\pm$1.31
Lee & Kim (2023) 90.10$\pm$0.18 83.66$\pm$0.46
+ SSNI-L 93.49$\pm$0.33 85.29$\pm$0.24
WRN-70-16 Nie et al. (2022) 92.90$\pm$0.40 82.94$\pm$1.13
+ SSNI-L 94.99$\pm$0.24 84.44$\pm$0.56
Wang et al. (2022) 93.10$\pm$0.51 85.03$\pm$0.49
+ SSNI-L 93.88$\pm$0.49 82.88$\pm$0.79
Lee & Kim (2023) 89.39$\pm$1.12 84.51$\pm$0.37
+ SSNI-L 92.64$\pm$0.40 84.90$\pm$0.09
@@ Appendix H Ablation Study on Score Norm
WRN-28-10 Nie et al. (2022) 89.71$\pm$0.72 47.98$\pm$0.64
+ SSNI-L 91.31$\pm$0.24 46.92$\pm$0.52
+ SSNI-N 92.84$\pm$0.18 47.20$\pm$1.22
Wang et al. (2022) 92.45$\pm$0.64 36.72$\pm$1.05
+ SSNI-L 93.15$\pm$0.92 35.72$\pm$1.33
+ SSNI-N 93.42$\pm$0.60 34.24$\pm$1.45
Lee & Kim (2023) 90.10$\pm$0.18 56.05$\pm$1.11
+ SSNI-L 93.40$\pm$0.49 54.52$\pm$0.46
+ SSNI-N 93.55$\pm$0.42 55.47$\pm$1.15

### v2

@@ 6 Limitation
Maximum level $t^{\rm S}$ for EPS. We use EPS to replace the single reference noise level with an integrated approach. Still, the maximum level $t^{\rm S}$ defining the upper bound of the expectation range is shared across samples, which may not be the optimal choice though. Refining sample-sensitive maximum levels could potentially improve the adaptability and robustness of EPS in different scenarios, and we leave it as future work.
The design of reweighting functions. The proposed reweighting functions (i.e., the linear and non-linear reweighting functions) may not be the optimal ones for SSNI. Further efforts are needed to explore data-driven or learning-based strategies to discover more effective function forms. Meanwhile, designing an effective reweighting function is an open question, and we leave it as future work.
Extra computational cost. The integration of an extra reweighting process will inevitably bring some extra cost. Further efforts are needed to optimize the trade-off between accuracy gains and computational overhead. Luckily, we find that this reweighting process is lightweight, making SSNI computationally feasible compared to existing DBP methods (see Section 5.6).
@@ Acknowledgements
JCZ is supported by the Melbourne Research Scholarship and would like to thank Shuhai Zhang and Huanran Chen for productive discussions. ZSY is supported by the Australian Research Council (ARC) with grant number DE240101089. FL is supported by the ARC with grant number DE240101089, LP240100101, DP230101540 and the NSF&CSIRO Responsible AI program with grant number 2303037. This research was supported by The University of Melbourne’s Research Computing Services and the Petascale Campus Initiative. Chaowei Xiao is supported by the Schmidt Sciences AI2050 fellow program.
@@ Appendix A Detailed Related Work
Adversarial attack. Adversarial examples (AEs) have emerged as a critical security concern in the development of AI systems in recent years (Szegedy et al., 2014; Goodfellow et al., 2015). These examples are typically generated by introducing imperceptible perturbations to clean inputs, which can cause a classifier to make incorrect predictions with high confidence. The methods used to generate such examples are known as adversarial attacks. One of the earliest attack methods, the fast gradient sign method (FGSM), which perturbs clean data in the direction of the gradient of the loss function (Goodfellow et al., 2015). Building on this idea, Madry et al. (2018) propose the projected gradient descent (PGD) attack, which applies iterative gradient-based updates with random initialization. AutoAttack (AA) (Croce and Hein, 2020) integrate multiple attack strategies into a single ensemble, making it a widely adopted benchmark for evaluating adversarial robustness. To address defenses that apply randomized input transformations, Athalye et al. (2018b) introduce the expectation over transformation (EOT) framework for computing more accurate gradients. Additionally, Athalye et al. (2018a) propose the backward pass differentiable approximation (BPDA), which approximates gradients using identity mappings to circumvent gradient obfuscation defenses. According to Lee and Kim (2023), the combination of PGD and EOT, i.e., PGD+EOT, is currently considered the most effective attack strategy against DBP methods. More recently, adversarial attacks that specifically designed for DBP methods are proposed. For example, Xue et al. (2023) propose diffusion-based projected gradient descent (Diff-PGD), which leverages an off-the-shelf diffusion model to guide perturbation optimization, enabling the generation of more stealthy AEs. Chen et al. (2024b) propose DiffAttack, which crafts perturbations in the latent space of diffusion models, rather than directly in pixel space, enabling the generation of human-insensitive yet semantically meaningful AEs through content-preserving structures.
@@
Adversarial defense. To counter the threat posed by adversarial attacks, numerous defense strategies have been developed, including adversarial detection (AD), adversarial training (AT), and adversarial purification (AP). (1) AD: AD is the most lightweight approach to defending against adversarial attacks is to detect and remove AEs from the input data. Earlier studies typically trained detectors tailored to specific classifiers or attack types, often overlooking the underlying data distribution, which limits their generalization to unseen attacks (Ma et al., 2018; Lee et al., 2018; Raghuram et al., 2021; Pang et al., 2022). Recently, statistical adversarial data detection has attracted growing attention for its ability to address this limitation. For instance, Gao et al. (2021) show that the maximum mean discrepancy (MMD) is sensitive to adversarial perturbations, and use distributional discrepancies between AEs and CEs to effectively identify and filter out AEs, even under previously unseen attacks. Building on this insight, Zhang et al. (2023) introduce a novel statistic called the expected perturbation score (EPS), which quantifies the average score of a sample after applying multiple perturbations. They then propose an EPS-based variant of MMD to capture the distributional differences between clean and adversarial examples more effectively. (2) AT: Vanilla AT (Madry et al., 2018) directly generates adversarial examples during training, encouraging the model to learn their underlying distribution. Beyond vanilla AT, various extensions have been proposed to improve its effectiveness. For instance, Zhang et al. (2019) introduce a surrogate loss optimized based on theoretical robustness bounds. Similarly, Wang et al. (2020) examine the role of misclassified examples in shaping model robustness and enhance adversarial risk via regularization techniques. From a reweighting perspective, Zhang et al. (2021) propose geometry-aware instance-reweighted AT (GAIRAT), which adjusts sample weights based on their distance to the decision boundary. Building on this, Wang et al. (2021) utilize probabilistic margins to reweight AEs in a continuous and path-independent manner. More recently, Zhang et al. (2024) suggest pixel-wise reweighting of AEs to explicitly direct attention toward critical image regions. (3) AP: AP typically employs generative models to transform AEs back into their clean counterparts before classification (Liao et al., 2018; Samangouei et al., 2018; Song et al., 2018; Naseer et al., 2020; Zhang et al., ). Within this context, diffusion-based purification (DBP) methods have emerged as a promising framework, exploiting the inherent denoising nature of diffusion models to filter out adversarial noise (Nie et al., 2022; Wang et al., 2022; Xiao et al., 2023; Lee and Kim, 2023). DBP works with pre-trained diffusion model by first corrupting an adversarial input through the forward process and then iteratively reconstructing it via the reverse process. This projects the input back onto the clean data manifold while stripping away adversarial artifacts. Wang et al. (2022) introduce input guidance during the reverse diffusion process to ensure the purified outputs stay close to the inputs. Lee and Kim (2023) propose a fine-tuned gradual noise scheduling for multi-step purifications. Bai et al. (2024) improve the reverse diffusion process by incorporating contrastive objectives.
@@
[…]
@@ C.1 DiffPure
Existing DBP methods generally follow the algorithm of DiffPure (Nie et al., 2022). DiffPure conducts evaluation on AutoAttack (Croce and Hein, 2020) and BPDA+EOT adaptive attack (Athalye et al., 2018a) to measure model robustness. DiffPure chooses optimal $t^{*}=100$ and $t^{*}=75$ on CIFAR-10 against threat models $\ell_{\infty}(\epsilon=8/255)$ and $\ell_{2}(\epsilon=0.5)$, respectively. It also tests on high-resolution dataset like ImageNet-1K with $t^{*}=150$ against threat models $\ell_{\infty}(\epsilon=4/255)$. Calculating exact full gradients of the defense system of DiffPure is impossible since one attack iteration requires 100 function calls (with $t^{*}=100$ and a step size of 1). DiffPure originally uses numerical SDE solver for calculating gradients. However, the adjoint method is insufficient to measure the model robustness since it relies on the performance of the underlying SDE solver (Zhuang et al., 2020; Lee and Kim, 2023; Chen et al., 2024a). Therefore, we apply surrogate process to efficiently compute gradients of direct back-propagation in our evaluation. To overcome memory constraint issue, we align the step size settings of denoising process in adversarial attack to 5 with the evaluation settings in (Lee and Kim, 2023) and keep the timestep $t^{*}$ consistent with DiffPure. For ImageNet-1K evaluation, we can only afford a maximum of 10 function calls for one attack iteration.
@@ C.2 GDMP
Lee and Kim (2023) evaluated GDMP with three types of guidance and concluded that No-Guidance provides the best robustness performance when using the surrogate process to compute the full gradient through direct backpropagation. In our evaluation, we incorporate the surrogate process with No-Guidance to evaluate GDMP. Since it is impossible to calculate the gradients of the full defense system, we use a surrogate process consisting of same number of purification steps but with larger step size in the attack (with 6 denoising steps and 10 denoising steps for PGD+EOT and BPDA+EOT attack, respectively). Notably, GDMP only uses one purification run with 45 forward steps to evaluate on ImageNet-1K, which we keep consistent with this setting.
@@ C.3 GNS
Lee and Kim (2023) emphasizes the importance of selecting optimal hyperparameters in DBP methods for achieving better robust performance. Hence, Lee and Kim (2023) proposed Gradual Noise-Scheduling (GNS) for multi-step purification, which is based on the idea of choosing the best hyperparameters for multiple purification steps. It is basically the same architecture as GDMP (no guidance), but with different purification steps, forward steps and denoising steps. Specifically, GNS sets different forward and reverse diffusion steps and gradually increases the noise level at each subsequent purification step. We just keep the same hyperparameter settings and also use an ensemble of 10 runs to evaluate the method.
@@ Appendix D Surrogate Process of Gradient Computation
The surrogate process is an efficient approach for computing approximate gradients through backpropagation, as proposed by (Lee and Kim, 2023). White-box adaptive attacks, such as PGD+EOT, involve an iterative optimization process that requires computing the exact full gradients of the entire system, result in high memory usage and increased computational time. DBP methods often include a diffusion model as an add-on purifier, which the model requires extensive function calls during reverse generative process. Hence, it is hard to compute the exact full gradient of DBP systems efficiently. The surrogate process takes advantage of the fact that, given a fixed total amount of noise, we can denoise it using fewer denoising steps (Song et al., 2021a), but the gradients obtained from the surrogate process slightly differ from the exact gradients. Instead of using the full denoising steps, we approximate the original denoising process with fewer function calls, which allows us to compute gradients by directly back-propagating through the forward and reverse processes.
There are other gradient computation methods such as adjoint method in DiffPure (Li et al., 2020; Nie et al., 2022). It leverages an underlying numerical SDE solver to solve the reverse-time SDE. The adjoint method can theoretically compute exact full gradient, but in practice, it relies on the performance of the numerical solver, which is insufficient to measure the model robustness (Zhuang et al., 2020; Lee and Kim, 2023; Chen et al., 2024a). Lee and Kim (2023) conducted a comprehensive evaluation of both gradient computation methods and concluded that utilizing the surrogate process for gradient computation poses a greater threat to model robustness. Hence, we use gradients obtained from a surrogate process in all our experiments.
@@ Appendix E Adaptive White-box PGD+EOT Attack for SSNI
0: clean data-label pairs (${\mathbf{x}},y)$; purifier $f_{p}$; classifier $f_{c}$; a noise level $T$; a score network $s_{\theta}({\mathbf{x}},T)$; objective function $\mathcal{L}$}; perturbation budget $\epsilon$; step size $\alpha$; PGD iterations $K$; EOT iterations NN.
@@ Appendix F Adaptive White-box BPDA+EOT Attack for SSNI
0: clean data-label pairs (${\mathbf{x}},y)$; purifier $f_{p}$; classifier $f_{c}$; approximation function $f_{\text{apx}}$; a noise level $T$; a score network $s_{\theta}({\mathbf{x}},T)$; objective function $\mathcal{L}$}; perturbation budget $\epsilon$; step size $\alpha$; PGD iterations $K$; EOT iterations NN.
@@ Appendix G Performance Evaluation of SSNI-L
[…]
@@
[…]
@@ Appendix H Ablation Study on Score Norm
[…]